OpenStudy (anonymous):

Let ABC be an equilateral triangle and let P be a point inside the triangle such that PA=3 PB=4 PC=5 Find the area of the triangle ABC.

OpenStudy (anonymous):

Area of ABC = \(\large \frac{x\sqrt3}{2}\cdot x \cdot \frac{1}{2}=\frac{x^2\sqrt3}{4} \)

OpenStudy (anonymous):

You do not know x.

OpenStudy (anonymous):

Hey, wouldnt the point P create 3 sub triangles within ABC? So you could calculate the total area by adding each sub area. Like, PA * PB over 2 + PB * PC over 2 + PA * PC over 2?

OpenStudy (anonymous):

still trying doc...:) law of cosines...,

OpenStudy (anonymous):

oh nevermind, thats just wrong

OpenStudy (anonymous):

\(\large x^2=3^2+4^2-2(3)(4)cos\theta \)

OpenStudy (anonymous):